A constant volume gas thermometer shows pressure reading 50cm and 90cm of mercury at 0 degree c and 100 degree c repectively .when the pressure reading is 60cm of mercury , the temperature is ????????????????????????????????????????????????????????
1
Expert's answer
2013-11-25T05:24:44-0500
if we suppose, that pressure of mercury depends linearly on temperature, than we can find how much does it changes when temperature rises on 1 degree:

(90 cm - 50 cm)/100 degree = 0.4 cm/1 degree.

hence, when changing on 60-50=10 cm:

h = 10/0.4 = 25 degree

And so, the temperature will be 25 c degree.